OpenStudy (anonymous):

sum inf k=1 (cos 1)^k diverge or converge if converge find the sum

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} (\cos 1)^k\] so nobody can solve this then

OpenStudy (anonymous):

I'm not sure if this is right, but I'll give it a shot. Since cos(1) = 0, the sum is an infinite addition of zeros. So maybe the answer is that it converges to zero?

OpenStudy (anonymous):

please do correct me if I'm wrong

OpenStudy (anonymous):

close

OpenStudy (anonymous):

true is converge but u need to use the geometry formula

OpenStudy (anonymous):

What is the geometry formula?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} ar^n-1= a+ar+ar^2+ar^3+... \] is converge if |r|<1 and its sum is \[\sum_{n=1}^{\infty} ar^n-1\]= a/1-r |r|<1 if |r|>=1 then is diverge

OpenStudy (anonymous):

this the formula